"Please help solve for M -7/8m = 1/4 am having trouble with fractions I am not working off text book am working off loose leaf sheet
Answer by checkley75(3666)

-7M/8=1/4 NOW CROSS MULTIPLY
-7M*4=8*1
-28M=8
M=8/-28
M=-8/28
M=-2/7
PROOF
(-7*-2/7)/8=1/4
(14/7)/8=1/4
2/8=1/4
1/4=1/4